Assume that Wal-Mart Stores, Inc. has decided to surface and maintain for 10 years a vacant lot next to one of its stores to serve as a parking lot for customers. Management is considering the following bids involving two different qualities of surfacing for a parking area of 11,700 square yards.

Bid A: A surface that costs $5 per square yard to install. This surface will have to be replaced at the end of 5 years. The annual maintenance cost on this surface is estimated at 25 cents per square yard for each year except the last year of its service. The replacement surface will be similar to the initial surface.

Bid B: A surface that costs $10 per square yard to install. This surface has a probable useful life of 10 years and will require annual maintenance in each year except the last year, at an estimated cost of 8 cents per square yard. (Use the tables below.)

Assume that Wal-Mart Stores, Inc. has de
Answer #1

Solution:

Surface Cost = Cost per square yard * Total square yards.

Annual maintenance cost = Annual maintainence per sqaure yard * Total square yards.

For Bid A-

Surface Cost = $5 * 11,700 = $58,500

Annual maintainence cost = $0.25 * 11,700 = $2,925.

As mentioned the 5th year and the 10th year will have no maintenance cost, while at the end of 5th year surface will be replaced.

For Bid B-

Surface Cost = $10 * 11,700 = $117,000

Annual maintenance cost = $0.08 * 11,700 = $936.

Present Value of Surface Cost = $117,000

Present Value of Annuity = Annuity * Annuity Factor

Present Value of Annuity = $936 * 6.24689 = $5,847.09

Present Value of Outflows for Bid B =  $117,000 + $5,847.09 = $122,847.09

Present Value of Outflows for Bid B =  $122,847.09.

Walmart should accept Bid A, since it has lower present value of outflows.